U.S. Steel will have 150,000 tons of steel to deliver in December 2019. The January 23, 2019 spot price of steel is $775 per ton. On Jan. 23, 2019 the company enters into a forward contract with General Motors for 150,000 tons of steel for delivery on December 1, 2019 with a forward price of $795 per ton. Suppose the steel spot price on Dec. 1, 2019 is $850 per ton. Answer q's about the following two scenarios.

A) Assuming that physical delivery of the steel actually occurs on the forward contract, how much revenue does U.S. Steel generate on the sale of steel to GM on Dec. 1, 2019?

B) Assume that U.S. Steel and GM instead agree to a cash settlement of the forward contract on Dec. 1, 2019. The cash settlement amount is calculated as the diff. between the forward price and the spot price times the quantity of steel. Further assume that U.S. steel sells 150,000 tons of steel to another counterparty in the spot market for the spot price on Dec. 1, 2019. Calculate the cash settlement amount on the forward contract, the revenue generated from the spot market transaction, and the total cash flow for U.S. Steel on December 1, 2019.
Answer #1

A) Total revenue generated by entering into forward contract is 1,50,000* 795 =11,92,50,000

B) Loss of entering into contract = 1,50,000*(795 - 850) = 82,50,000, as the company is doing sales the amount received will be less by entering into forward contract.

Further sale in the spot market = 1,50,000 * 850 = 12,75,00,000

Therefore, total cash flows = 12,75,00,000 - 82,50,000 = 11,92,50,000